# What You Say = What You Want? Teaching Humans to Articulate Requirements for LLMs

### v1

@@ 7.2. Reflection on the ROPE Paradigm
Along with prior work on requirement-oriented evaluation (Kim et al., 2023), we believe that ROPE moves us towards a future where requirements serve as the central interface between AI and humans. Here, we discuss key skills humans and LLMs need to develop to prepare for a ROPE future.
@@
To ensure success of ROPE, multiple skills need to be cultivated among humans. Our training successfully helped users develop the ability to extract requirements from given examples and express them completely and correctly, but our study also revealed more opportunities in requirement iteration and testing skills of an end-to-end prompting procedure. Specifically, users need to identify examples to compare LLM outputs and adjust the requirements accordingly.
For example, users should create diverse and representative examples to identify and test requirements, particularly for ambiguous or open-ended tasks involving edge cases or complex interactions. In our study, many users, especially those unfamiliar with games like Connect4, missed the testing scenarios like a tie game, which led them to omit the requirement ‘‘Display a "Tie Game!" message if no player wins’’ in their prompt. This is consistent to findings in debugging training (Ma et al., 2024) and other end-user prompt engineering work (Zamfirescu-Pereira et al., 2023), as novices often create too few test examples. Future studies should support users to generate more testing examples.
Additionally, users should iterate requirement granularity based on task “LLM-hardness.” This involves observing LLM failures and refining requirements, essentially developing a theory of mind for LLMs (Wang et al., 2024). Some of our participants was already thinking about “what is obvious enough for ChatGPT to already understand/perform well on” (S28) and adjust their prompt accordingly (although not necessarily correctly). In our training, we introduced the concept of requirement specificity in Tetris, where users generate requirements in two levels of granularity (i.e., main steps and details, Section 4.3). However, some tasks required even greater specificity, such as generating non-standard Connect4 gameplay elements ‘‘Cross out winning cells’’. This ability to adjust specificity in requirements can be influenced by domain expertise, for example, expert developers can iterate prompts with more accurate details to successfully generate code while students often struggle (Nam et al., 2023). Future studies should offer adaptive support for adjusting requirement granularity based on task complexity and user expertise.
@@
Second, optimizers can also assist test data synthesis. More advanced optimizers need to iterate prompts on validation datasets and automatic objective functions, similar to machine learning. Insights from software engineering test automation (Just et al., 2014) or NLP approaches to extract or synthesize test cases from natural language requirements could be helpful (Garousi et al., 2020; Yang et al., 2023; Yu et al., 2024; Yang et al., 2024b).
Finally, LLMs or optimizers should explicitly communicate their interpretations and reveal biases in their implementation. As we observed in our study, optimizers can misinterpret requirements or add unnecessary details (e.g., “interactive game” can lead to unwanted features like Flask integration). LLMs should explicitly communicate their understandings of user intent, such as translating ambiguous natural language requirements into more formal specifications (Liu et al., 2023b; Ym et al., 2023), asking clarifying questions (Qian et al., 2024; Lahiri et al., 2022; Chen et al., 2024), or extracting and refining generated criteria (Kim et al., 2023; Yuan et al., 2024). This is not a perfect solution, as models’ translations may still reflect their biases and models may miss to ask about LLM-hard requirements. Nonetheless, to correctly implement requirements in LLM outputs when what is LLM-hard keep evolving, humans need to learn when and how to specify requirements with more granularity, and models need to keep communicating what requirements might be LLM-hard or LLM-conflict.
@@ 8. Conclusion
In this work, we advocate for focusing prompt engineering effort on human-centered tasks, ensuring users include all necessary requirements in their prompt to achieve goals. We introduce the Requirement-Oriented Prompt Engineering (ROPE) paradigm, and design training materials where prompting novices practice requirement articulation on complex prompting tasks. We also develop aligned assessment metrics capturing both intrinsic quality of user prompts in terms of requirement quality, as well as the extrinsic quality of prompt effectiveness in achieving intended outcomes. By providing targeted feedback on requirements, our ROPE system helps users produce higher-quality requirements and prompts more effectively than traditional prompt engineering training. As we look to the future, it is clear that the demand for LLM-based applications will only grow. As LLMs become more integrated into complex task-solving for more users, the ability to clearly articulate requirements will be key to effectively guide LLMs. We believe that users should be equipped with the foundational skills to prepare for the ROPE future, and the right training will empower users to harness the full potential of LLMs.
@@ Acknowledgements.
Thanks to all the participants for the pilot, interview, and user study in this work. Thanks to Kelly Rivers, Michael Taylor, Michael Hilton, Michael Xieyang Liu, Xinran Zhao, Chenyang Yang, Lauren Sands, Austin Schick, David Kosbie, and Daniel Anderson for all the insights, advice, and help. Thanks to CMU HCII faculty and Ken’s lab for feedback.

### v2

@@ 7.2. Reflection on the ROPE Paradigm
Along with prior work on requirement-oriented evaluation (Kim et al., 2024), we believe that ROPE moves us towards a future where requirements serve as the central interface between AI and humans. Here, we discuss key skills humans and LLMs need to develop to prepare for a ROPE future. These insights can also be useful for designing better prompting support tools.
@@
To ensure the success of ROPE, multiple skills need to be cultivated among humans. Our training successfully helped users develop the ability to extract requirements from given examples and express them completely and correctly, but our study also revealed more opportunities in requirement iteration and testing skills of an end-to-end prompting procedure. Specifically, users need to identify examples to compare LLM outputs and adjust the requirements accordingly.
For example, users should create diverse and representative examples to identify and test requirements, particularly for ambiguous or open-ended tasks involving edge cases or complex interactions. In our study, many users, especially those unfamiliar with games like Connect4, missed the testing scenarios like a tie game, which led them to omit the requirement ‘‘Display a "Tie Game!" message if no player wins’’ in their prompt. This is consistent to findings in debugging training (Ma et al., 2024) and other end-user prompt engineering work (Zamfirescu-Pereira et al., 2023), as novices often create too few test examples. Future studies should support users to generate more nuanced testing examples, supporting comprehensive hypothesis testing in prompt engineering (Arawjo et al., 2024).
Additionally, users should iterate requirement granularity based on task “LLM-hardness.” This involves observing LLM failures and refining requirements, essentially developing a theory of mind for LLMs (Wang et al., 2024b). Some of our participants were already thinking about “what is obvious enough for ChatGPT to already understand/perform well on” (S28) and adjust their prompt accordingly (although not necessarily correctly). In our training, we introduced the concept of requirement specificity in Tetris, where users generate requirements in two levels of granularity (i.e., main steps and details, Section 4.3). However, some tasks required even greater specificity, such as generating non-standard Connect4 gameplay elements ‘‘Cross out winning cells’’. This ability to adjust specificity in requirements can be influenced by domain expertise, for example, expert developers can iterate prompts with more accurate details to successfully generate code while students often struggle (Nam et al., 2023). Future studies should offer adaptive support for adjusting requirement granularity based on task complexity and user expertise.
@@
Second, optimizers can also assist test data synthesis. More advanced optimizers need to iterate prompts on validation datasets and automatic objective functions, similar to machine learning. Insights from software engineering test automation or NLP approaches to extract or synthesize test cases from natural language requirements could be helpful (Garousi et al., 2020; Yang et al., 2023; Yu et al., 2024; Yang et al., 2024b).
Finally, LLMs or optimizers should explicitly communicate their interpretations and reveal biases in their implementation. As we observed in our study, optimizers can misinterpret requirements or add unnecessary details (e.g., “interactive game” can lead to unwanted features like Flask integration). LLMs should explicitly communicate their understandings of user intent, such as translating ambiguous natural language requirements into more formal specifications (Liu et al., 2023b; Lee et al., 2022), asking clarifying questions (Qian et al., 2024; Lahiri et al., 2022; Chen et al., 2024), or extracting and refining generated criteria (Kim et al., 2024; Yuan et al., 2024; Wang et al., 2024a). This is not a perfect solution, as models’ translations may still reflect their biases and models may miss to ask about LLM-hard requirements. Nonetheless, to correctly implement requirements in LLM outputs when what is LLM-hard keeps evolving, humans need to learn when and how to specify requirements with more granularity, and models need to keep communicating what requirements might be LLM-hard or LLM-conflict.
@@ 8. Conclusion
In this work, we advocate for focusing prompt engineering effort on human-centered tasks, ensuring users include all necessary requirements in their prompt to achieve goals. We introduce the Requirement-Oriented Prompt Engineering (ROPE) paradigm, and design training materials where prompting novices practice requirement articulation on complex prompting tasks. We also develop aligned assessment metrics capturing both the intrinsic quality of user prompts in terms of requirement quality, as well as the extrinsic quality of prompt effectiveness in achieving intended outcomes. By providing targeted feedback on requirements, our ROPE system helps users produce higher-quality requirements and prompts more effectively than traditional prompt engineering training. As we look to the future, it is clear that the demand for LLM-based applications will only grow. As LLMs become more integrated into complex task-solving for more users, the ability to clearly articulate requirements will be key to effectively guiding LLMs. We believe that users should be equipped with the foundational skills to prepare for the ROPE future, and the right training will empower users to harness the full potential of LLMs.
@@ Acknowledgements.
Thanks to all the participants for the pilot, interview, and user study in this work. Thanks to Kelly Rivers, Michael Taylor, Michael Hilton, Michael Xieyang Liu, Xinran Zhao, Lauren Sands, Austin Schick, David Kosbie, and Daniel Anderson for all the insights, advice, and help. Thanks to the CMU HCII faculty and Ken’s lab for feedback. Thanks to the National Science Foundation (award CNS-2213791, 2414915) and gift fund from Google for support of this work. Thanks to the OpenAI research credit program and Amazon AI research gift fund.